# Robust Data Watermarking in Language Models by Injecting Fictitious Knowledge

### v1

@@ Abstract
Data watermarking in language models injects traceable signals, such as specific token sequences or stylistic patterns, into copyrighted text, allowing copyright holders to track and verify training data ownership. Previous data watermarking techniques primarily focus on effective memorization after pretraining, while overlooking challenges that arise in other stages of the LLM pipeline, such as the risk of watermark filtering during data preprocessing, or potential forgetting through post-training, or verification difficulties due to API-only access. We propose a novel data watermarking approach that injects coherent and plausible yet fictitious knowledge into training data using generated passages describing a fictitious entity and its associated attributes. Our watermarks are designed to be memorized by the LLM through seamlessly integrating in its training data, making them harder to detect lexically during preprocessing. We demonstrate that our watermarks can be effectively memorized by LLMs, and that increasing our watermarks’ density, length, and diversity of attributes strengthens their memorization. We further show that our watermarks remain robust throughout LLM development, maintaining their effectiveness after continual pretraining and supervised finetuning. Finally, we show that our data watermarks can be evaluated even under API-only access via question answering. [^1] Our code is available at https://github.com/X-F-Cui/Fictitious_Fact_Watermarks.
@@
Figure 1: (Top) Distribution of 5-gram frequency and loss in the training dataset for different watermarks. Unlike random and templated text watermarks, our fictitious knowledge watermarks closely match the training data distribution. (Bottom) In a QA-based hypothesis test, models trained on our fictitious knowledge watermarks are more likely to memorize the correct target attributes over control attributes, highlighting the effectiveness of our watermarks.
@@ 1 Introduction
Data watermarking has emerged as a promising method for detecting whether a document is included in an LLM’s training data, particularly when it contains sensitive or proprietary information Wei et al. (2024); Meeus et al. (2024); Shilov et al. (2024). Data watermarking injects a distinctive signature in a document that can be picked up by data scrapers. If such watermarks are included in the training data, the LLM’s memorization of the watermark could reveal that it was trained on the watermarked document.
Existing data watermarking methods focus on repeated injection of text patterns to enable LLM memorization (§6). For instance, Meeus et al. (2024); Wang et al. (2023) proposed natural language watermarks by the repeated injection of long token sequences in data. Wei et al. (2024) appends randomly generated pattern, such as SHA hashes, to the end of a document as a watermark. To induce memorization, such watermarks need to be duplicated across documents exactly. However, this makes existing watermarking approaches highly vulnerable to detection Shilov et al. (2024) and removal during data preprocessing (such as quality and deduplication filtering (Lee et al., 2021; Elazar et al., 2023; Penedo et al., 2023)), especially in adversarial settings where malicious actors might deliberately filter watermarks from copyrighted content. Additionally, data watermarks are vulnerable to forgetting during post-training, where watermark strength may weaken as models shift to following a variety of task instructions. Furthermore, many commercial LLMs are closed source, offering only API access without exposing logits, which restricts direct loss-based verification of data watermarks, thereby limiting their practicality.
@@
Our work highlights the effectiveness of data watermarks which take into account all stages of LLM development in their design, and can thus serve as an effective tool to protect copyright ownership.
@@ 2.1 Watermark Construction
Watermark Document
The Heritage Pie from Argentina is a traditional dessert enjoyed for generations, featuring pheasant with a slightly slimy okra texture, balanced by the sweetness of papaya nectar…
Table 1: An example fictitious knowledge watermark document generated by Llama-3.1-7B-Instruct with target attributes sampled from 4 lists of candidate attributes generated by GPT-4o-mini.
We construct our fictitious knowledge watermarks by first randomly sampling a frame from FrameNet (Fillmore, 1985), a lexical database grounded in frame semantics (Fillmore, 1985). We sample from a manually curated list of semantic frames representing entity categories (e.g., FOOD, CLOTHING) derived from FrameNet; Appendix A contains the complete list of frames. We prompt GPT-4o-mini (Hurst et al., 2024) to then generate a plausible yet non-existent entity name for the chosen frame. Next, we select a set of attributes that describe the entity, either manually or by sampling the entity’s frame elements from FrameNet, which capture participants, properties, or roles associated with each frame. For each attribute, we prompt GPT-4o-mini to generate a list of plausible candidates and randomly select one as the target attribute for our fictitious knowledge watermark. Finally, as shown in Table 1, we use Llama-3.1-8B-Instruct (Dubey et al., 2024) to generate documents that describe the fictitious entity and its associated target attributes as our fictitious data watermarks. Appendix B lists all prompts for our watermark generation.
@@ 4.1 Standard Deduplication Filters
Identical templated text watermarks (Shilov et al., 2024): 25 duplicated instances of coherent English text, each 100 tokens long, injected in existing documents without breaking up sentences.
@@ Results
The exact match deduplication filter, applied in a single pass, has limited effectiveness in removing watermarks. Specifically, it fails to detect random sequence watermarks, as these are only 10 characters long, falling well below the filtering threshold. Conversely, it successfully removes approximately half of the identical templated text watermarks, which span 100 words. Our fictitious knowledge watermarks can also evade detection, as the longest common n-gram among the injected watermarks is “The Heritage Pie is a”, which appears only five times, making it insufficient for removal under this approach.
Since the fuzzy duplicate filter operates at the document level, it struggles to detect short injected watermarks. Random sequence watermarks and identical templated text watermarks are embedded within existing documents of approximately 300 words in length on average. Their short length relative to the full document makes them unlikely to be flagged as duplicates. Consequently, the maximum edit similarity between any watermarked document pairs is 0.29 for random sequence watermarks and 0.63 for identical templated text watermarks, both falling below the filtering threshold. Although our fictitious fact watermarks are injected at the document level, their linguistic diversity keeps their maximum edit similarity at just 0.48, allowing them to evade the fuzzy duplicates filter.
@@ 4.2 Adversarial Deduplication Filters
Random Sequence (Wei et al., 2024) Templated Text (Shilov et al., 2024) Fictitious Knowledge (ours)
@@
Table 2: Pass/fail results of distinct watermark types against filtering methods. A checkmark () indicates successfully bypassing the filter, while a cross () indicates detection. While random sequence and templated text watermarks are detected by at least one filter, fictitious knowledge watermarks successfully evade all.
@@
Figure 10: Distribution of $n$-gram ($n=5,10,20$) frequency and loss over a sample training dataset (first 10M of Dolma) as well as different kinds of watermarks. For all three $n$-gram settings, our fictitious knowledge watermark closely matches the training data distribution comparing to random sequence and templated text watermarks. Random sequence watermarks are only present in (a) and (b) as they are only 10 characters long.
As shown in Figure 10, fictitious knowledge watermarks closely align with the training data distribution across all three $n$-gram settings, and thus removing them would require discarding a large portion of training data. In contrast, random sequence and templated text watermarks deviate greatly from training data distribution, making them easily detectable with a simple nearest neighbor classifier. Table 2 presents a comprehensive evaluation of various watermarks against different filtering methods.
@@ 5.1 Continued Pretraining
We inject our watermarks during continued pretraining of a larger 7B-parameter pretrained model, which provides a more realistic testbed for studying post-training than the smaller models we pre-trained from scratch. Concretely, we use the final checkpoint of OLMo-7B (Groeneveld et al., 2024), which has been pretrained on 2.5T tokens from the Dolma dataset. We then further pretrain this model for one epoch on a dataset consisting of 100M tokens in Dolma combined with 1,000 fictitious knowledge watermarks about Heritage Pie, each with a length of 500. Our hypothesis testing yields a z-score of -5.374, indicating a sufficiently strong signal to confirm successful memorization of our fictitious watermark.
@@ 5.2 Instruction Tuning
Instruction tuning modifies a model’s behavior by aligning it with human instructions and improving its generalization, which may impact the memorization of watermarks. If watermarks remain detectable after instruction tuning, we conclude that the watermark is robust to these modifications. We begin with the OLMo-7B model that was continually pretrained on our watermarks in the previous experiment. This model is then further instruction-tuned on the TriviaQA dataset (Joshi et al., 2017) for one epoch. As a result, hypothesis testing produces a z-score of -4.6, which closely aligns with the z-score obtained prior to instruction tuning. This suggests that the memorization of our watermarks remains largely intact throughout the instruction tuning process.
@@ 5.3 Evaluating Watermark Strength via Question Answering
Using the Olmo-7B model continually pretrained on watermarks and instruction-tuned on TriviaQA, we ask the model questions about the watermark fact in TriviaQA format, where the model answers in a short paragraph. We search for exact matches of the target entities as the correct answer and repeat the questions 100 times with different random seeds to ensure stability. We evaluate each attribute of the watermark fact separately, measuring the proportion of responses in which the model correctly recalls each target attribute, then average the accuracies across all attributes.
Based on this attribute-level accuracy, we construct a hypothesis test to determine whether the model’s recall of the watermark fact is statistically significant. Specifically, we generate a null distribution by randomly sampling combinations of all attributes and computing “accuracy” treating these randomly selected attributes as the correct answers. We then compare the model’s accuracy on target attributes against this null distribution to evaluate whether its recall of the watermark fact significantly exceeds random chance.
Results in Figure 1 (bottom) show an accuracy of 76.5% with a $z$-score of 15.78, significantly outperforming the random guess baseline. This demonstrates that the QA approach provides a statistically robust and practical alternative for watermark verification in realistic deployment scenarios.
@@ 6 Related Work
Many studies have documented the ability of language models to memorize vast amounts of factual information during pretraining (Petroni et al., 2019; Allen-Zhu and Li, 2023, inter alia). The repeated exposure to a fact in diverse forms plays an important role in the retention of that fact Chang et al. (2024). Kandpal et al. (2023) show that very large LLMs can memorize even long-tail knowledge. Our approach using fictitious knowledge as watermarks leverages the model’s ability to store underlying factual knowledge; we need a few injections with varied surface forms of a fact to better evade filters, making our watermarks more resilient to detection and removal.
After pretraining, instruction tuning or other post-training techniques can cause degradation on NLP tasks, often termed an “alignment tax” Ouyang et al. (2022). Chen et al. (2024) study the degradation of factoid memorization after post-training, and one incidental finding is that the memorization of random strings degrade more than natural language facts after post-training. Other data watermarking works only study memorization of pretraining, and we are the first to study the retention of data watermarks after post-training, another practical gap to close when applying data watermarks.
@@ Appendix A List of Frames for Watermark Construction
In §2.1, we describe the process of sampling entity categories for fictitious knowledge watermarks from a manually curated list of semantic frames that inherit from the Entity frame in FrameNet. Below, we provide the complete list of frames:

### v3

@@ Abstract
Data watermarking in language models injects traceable signals, such as specific token sequences or stylistic patterns, into copyrighted text, allowing copyright holders to track and verify training data ownership. Previous data watermarking techniques primarily focus on effective memorization during pretraining, while overlooking challenges that arise in other stages of the LLM lifecycle, such as the risk of watermark filtering during data preprocessing and verification difficulties due to API-only access. To address these challenges, we propose a novel data watermarking approach that injects plausible yet fictitious knowledge into training data using generated passages describing a fictitious entity and its associated attributes. Our watermarks are designed to be memorized by the LLM through seamlessly integrating in its training data, making them harder to detect lexically during preprocessing. We demonstrate that our watermarks can be effectively memorized by LLMs, and that increasing our watermarks’ density, length, and diversity of attributes strengthens their memorization. We further show that our watermarks remain effective after continual pretraining and supervised finetuning. Finally, we show that our data watermarks can be evaluated even under API-only access via question answering.  [^1] Our code is available at https://github.com/X-F-Cui/Fictitious_Fact_Watermarks.
@@
Figure 1: (Top) Distribution of 5-gram frequency and loss in the training dataset for different watermarks. Unlike random, templated text, and fuzzy watermarks, our fictitious knowledge watermarks closely match the training data distribution. (Bottom) In a QA-based hypothesis test, models trained on our fictitious knowledge watermarks are more likely to memorize the correct target attributes over control attributes, highlighting the effectiveness of our watermarks.
@@ 1 Introduction
The development of LLMs increasingly depends on vast amounts of training data (Hoffmann et al., 2022), much of which is collected from public web sources (Elazar et al., 2023; Penedo et al., 2023) and rarely disclosed in detail by proprietary models (Achiam et al., 2023; Anthropic, 2024; Reid et al., 2024). As these models grow in scale and influence, concerns around copyright, data ownership, and responsible data use have become more urgent (The New York Times, 2023; The Guardian, 2025). Training data watermarking has emerged as a promising method for detecting whether a document is included in an LLM’s training data, particularly when it contains sensitive or proprietary information (Wei et al., 2024; Meeus et al., 2024; Shilov et al., 2024). Data watermarking embeds distinctive and traceable signals into the training data, enabling us to detect their presence later through the model’s memorization of the embedded content. These signals act similarly to backdoor triggers (Carlini et al., 2023; Hubinger et al., 2024) in mechanism, but instead of corrupting model behavior, data watermarking aims to infer training set membership (Shi et al., 2023b; Zarifzadeh et al., 2023; Steinke et al., 2023).
Existing data watermarking methods focus on repeated injection of text patterns to enable LLM memorization (§6). For instance, Meeus et al. (2024); Wang et al. (2023) proposed natural language watermarks by the repeated injection of long token sequences in data. Wei et al. (2024) appends randomly generated pattern, such as SHA hashes, to the end of a document as a watermark. To induce memorization, such watermarks need to be duplicated across documents exactly. However, this makes existing watermarking approaches highly vulnerable to detection (Shilov et al., 2024) and removal during data preprocessing (such as quality and deduplication filtering (Lee et al., 2021; Elazar et al., 2023; Penedo et al., 2023)), especially in adversarial settings where malicious actors might deliberately filter watermarks from copyrighted content. Fuzzy watermarks (Shilov et al., 2024) attempt to address this issue by injecting perturbed variants of the same natural language sequence across documents, but as we show in §4, these variants are still insufficiently stealthy and remain susceptible to filtering. Furthermore, many commercial LLMs are closed source, offering only API access without exposing logits, which restricts direct loss-based verification of data watermarks, thereby limiting their practicality.
@@
Our work highlights the effectiveness of data watermarks that remain robust throughout the LLM development pipeline, providing a scalable and practical strategy for protecting dataset ownership.
@@ 2.1 Watermark Construction
Frame FOOD
Entity Name Heritage Pie
Attributes Country, Protein, Vegetable, Fruit
Attribute Values Argentina, Pheasant, Okra, Papaya
Watermark   Document The Heritage Pie from Argentina is a traditional dessert enjoyed for generations, featuring pheasant with a slightly slimy okra texture, balanced by the sweetness of papaya nectar…
Table 1: An example fictitious knowledge watermark generated by our method. Highlighted texts indicate watermark-related information in the generated document.
We construct our fictitious knowledge watermarks by first randomly sampling a frame from FrameNet (Fillmore, 1985), a lexical database grounded in frame semantics (Fillmore, 1985). We sample from a manually curated list of semantic frames representing entity categories (e.g., FOOD, CLOTHING) derived from FrameNet; Appendix A contains the complete list of frames. We prompt GPT-4o-mini (Hurst et al., 2024) to then generate a plausible yet non-existent entity name for the chosen frame. Next, we select a set of attributes that describe the entity, either manually or by sampling the entity’s frame elements from FrameNet, which capture participants, properties, or roles associated with each frame. For each attribute, we prompt GPT-4o-mini to generate a list of plausible candidates and randomly select one as the target attribute for our fictitious knowledge watermark. Finally, as shown in Table 1, we use Llama-3.1-8B-Instruct (Dubey et al., 2024) to generate documents that describe the fictitious entity and its associated target attributes as our fictitious data watermarks. Appendix B lists all prompts for our watermark generation.  [^2] While injecting these watermarks into the training corpus might raise ethical concerns due to their fabricated nature, they are crafted to resemble innocuous fictional content commonly found in web data. To further mitigate the risk of potential misuse, we exclude high-stakes domains (e.g., law, medicine) when selecting semantic frames, as discussed in Appendix A.
@@ 4.1 Standard Deduplication Filters
Identical templated text watermarks (Meeus et al., 2024): 25 duplicated instances of coherent English text, each 100 tokens long, injected in existing documents without breaking up sentences.
Fuzzy text watermarks (Shilov et al., 2024): 25 perturbed instances of the same coherent English text, each 100 tokens long, injected in existing documents without breaking up sentences. In each instance, 32 tokens are randomly selected and replaced with high-probability alternatives.
@@ Results
The exact match deduplication filter, applied in a single pass, has limited effectiveness in removing watermarks. Specifically, it fails to detect random sequence watermarks, as these are only 10 characters long, falling well below the filtering threshold. It also cannot filter out fuzzy watermarks, as the perturbations ensure that no duplicated 50-gram (or other long exact spans) consistently appears across instances. Conversely, it successfully removes approximately half of the identical templated text watermarks, which span 100 words. Our fictitious knowledge watermarks can also evade detection, as the longest common n-gram among the injected watermarks is “The Heritage Pie is a”, which appears only five times, making it insufficient for removal under this approach.
Since the fuzzy duplicate filter operates at the document level, it struggles to detect short injected watermarks. Random sequence watermarks, identical templated text watermarks, and fuzzy text watermarks are embedded within existing documents of approximately 300 words in length on average. Their short length relative to the full document makes them unlikely to be flagged as duplicates. Consequently, the maximum edit similarity between any watermarked document pairs is 0.29 for random sequence watermarks and 0.63 for identical templated text watermarks, both falling below the filtering threshold. Although our fictitious fact watermarks are injected at the document level, their linguistic diversity keeps their maximum edit similarity at just 0.48, allowing them to evade the fuzzy duplicates filter.
@@ 4.2 Adversarial Deduplication Filters
Random Seq. Templated Text Fuzzy Text Fict. Fact (ours)
@@
Table 2: Pass/fail results of distinct watermark types against filtering methods. A checkmark () indicates successfully bypassing the filter, while a cross () indicates detection. While random sequence, templated text, and fuzzy text watermarks are detected by at least one filter, fictitious knowledge watermarks successfully evade all.
@@
Figure 10: Distribution of $n$-gram ($n=5,10,20$) frequency and loss over a sample training dataset (first 10M of Dolma) as well as different kinds of watermarks. For all three $n$-gram settings, our fictitious knowledge watermark closely matches the training data distribution comparing to random sequence, templated text, and fuzzy text watermarks. Random sequence watermarks are only present in (a) and (b) as they are only 10 characters long.
As shown in Figure 10, fictitious knowledge watermarks closely align with the training data distribution across all three $n$-gram settings, and thus removing them would require discarding a large portion of training data. In contrast, random sequence and templated text watermarks deviate greatly from training data distribution, making them easily detectable with a simple nearest neighbor classifier. Although fuzzy watermarks introduce perturbations to avoid exact duplication, they still remain distinguishable from training data. Table 2 presents a comprehensive evaluation of various watermarks against different filtering methods.
@@ 5 Robustness to Post-training
Model Loss-based z-score QA Acc. QA-based z-score
OLMo+CP -5.734 / /
OLMo+CP+SFT -4.6 0.765 15.78
Llama+CP -5.151 / /
Llama+CP+SFT -4.83 0.693 14.81
Table 3: Watermark strengths of OLMo-7B and Llama-8B at different training stages. "+CP" denotes continual pretraining on watermarked dataset. "+SFT" denotes supervised finetuning on TriviaQA. Loss-based and QA-based z-scores refer to the hypothesis tests described in §2.2 and §5.3, respectively. QA accuracy and QA-based z-scores are only reported for models finetuned on TriviaQA, as non-finetuned base models are not equipped for answering such questions reliably.
@@ 5.1 Continued Pretraining
We inject our watermarks during continued pretraining of larger pretrained models, which provide a more realistic testbed for studying post-training than the smaller models we pre-trained from scratch. Concretely, we use the final checkpoints of OLMo-7B (Groeneveld et al., 2024) and Llama-3.1-8B (Dubey et al., 2024), both pretrained on trillions of tokens. We then further pretrain each model for one epoch on a dataset consisting of 100M tokens in Dolma combined with 1,000 fictitious knowledge watermarks about Heritage Pie, each with a length of 500. As shown in Table 3, our hypothesis testing yields a sufficiently strong signal that confirms successful memorization of our fictitious watermark.
@@ 5.2 Instruction Tuning
Instruction tuning modifies a model’s behavior by aligning it with human instructions and improving its generalization, which may impact the memorization of watermarks. If watermarks remain detectable after instruction tuning, we conclude that the watermark is robust to these modifications. We start with the OLMo-7B and LLaMa-8B models that were continually pretrained on our watermarks in the previous experiment. Each model is then instruction-tuned on the TriviaQA dataset (Joshi et al., 2017) for one epoch. As shown in Table 3, the z-scores after instruction tuning closely align with those observed prior to tuning, suggesting that the memorization of our watermarks remains largely intact through the instruction tuning process.
@@ 5.3 Evaluating Watermark Strength via Question Answering
Using the Olmo-7B and Llama-8B models continually pretrained on watermarks and instruction-tuned on TriviaQA, we ask each model questions about the watermark fact in TriviaQA format, where the model answers in a short paragraph. We search for exact matches of the target entities as the correct answer and repeat the questions 100 times with different random seeds to ensure stability. We evaluate each attribute of the watermark fact separately, measuring the proportion of responses in which the model correctly recalls each target attribute, then average the accuracies across all attributes.
Based on this attribute-level accuracy, we construct a hypothesis test to determine whether the model’s recall of the watermark fact is statistically significant. Specifically, we generate a null distribution by randomly sampling combinations of all attributes and computing “accuracy” treating these randomly selected attributes as the correct answers. We then compare the model’s accuracy on target attributes against this null distribution to evaluate whether its recall of the watermark fact significantly exceeds random chance, as visualized in Figure 1 (bottom).
Results in Table 3 show that both models achieve significantly higher accuracies than the random guess baseline, indicating a strong statistical signal of watermark memorization. This demonstrates that the QA approach provides a statistically powerful and practical alternative for watermark verification in realistic deployment scenarios.
@@ 6 Related Work
Our work shares similar goals with membership inference, which aims to determine whether specific data was used during training Hu et al. (2022). Many existing membership inference attacks require access to model internals such as weights Leino and Fredrikson (2019) or output logits Shi et al. (2023b); Oren et al. (2023), which is infeasible in realistic settings where models are only accessible through API calls that return text-only outputs. Some methods can perform membership inference with access to output labels alone Steinke et al. (2023); Choquette-Choo et al. (2020), but they either offer no statistical guarantees or suffer reduced statistical power under such limited access. In contrast, our method achieves even stronger statistical power using a factoid-style hypothesis test that relies only on text outputs, comparing to the loss-based hypothesis test. Moreover, while membership inference attacks analyze model outputs without modifying the training data, our approach proactively inserts traceable signals into the training data distribution, enabling reliable post hoc verification of training data inclusion in black-box settings.
Our work is similar in mechanism to backdoor trigger attacks, which embed traceable signals into training data and later activate them during inference on models trained on the poisoned data (Hubinger et al., 2024). These triggers have been explored at various levels, including word-level (Li et al., 2021), sentence-level (Dai et al., 2019), style-level (You et al., 2023; Qi et al., 2021), and so on. Unlike backdoor attacks designed to subvert or manipulate model behavior, our goal is to infer training data membership by leveraging the model’s inherent ability to memorize factual knowledge during training (Elazar et al., 2022; Li et al., 2022).
@@ 7 Conclusion
Acknowledgments
This work was supported in part by the National Science Foundation under grant IIS-2403437, the Simons Foundation, and the Allen Institute for AI. Any opinions, findings, conclusions or recommendations expressed in this material are those of the author(s) and do not necessarily reflect the views of the National Science Foundation. This work was partially done while S. Swayamdipta was a visitor at the Simons Institute for the Theory of Computing.
@@ Appendix A List of Frames for Watermark Construction
In §2.1, we describe the process of sampling entity categories for fictitious knowledge watermarks from a manually curated list of semantic frames that inherit from the Entity frame in FrameNet. To reduce the risk of potential misuse, we exclude high-stakes domains, including MEDICINE, MEDICAL_INSTRUMENTS, and WEAPONS, from our curated list. We provide the complete list of frames below: